I tried Krytox around the O-ring and also tightening the screws around the valve. The leaking persists at roughly the same rate.
I tried pumping down the JPL PD chamber to test the new PD at cryo temperatures. Unfortunately, the chamber can;t get past about 6E-3 Torr with the pump on. As soon as I turned off the pump the pressure rose to around 2 Torr over 20 minutes or so.
I extricated the chamber from the pedestals, flipped it and removed the bottom plate. I cleaned the O-ring with isopropanol and wiped down the mating surface on the chamber (also with iso). I replaced the plate and tightened the screws. Then I returned the chamber to the table and reconnected it to the vacuum system. I tried pumping down once again but I saw pretty much exactly the same situation as before (pressure bottoming out around 6E-3 Torr and then rising quickly again when the pump was turned off).
I guess it's possible that the O-ring is damaged - although I couldn't see anything obivous. We didn't mess around with the viewport (when we replaced the diode a few weeks ago) so I'm hoping there is no issue there.